Fermi National Laboratory


Accelerator Update for April 13, 2005 - April 15, 2005

Wednesday April 13
The day shift began with Operations monitoring store 4097, with the Antiproton Source (Pbar) stacking , and with the MiniBooNE and Meson MCenter experiments taking beam.

A Tevatron (TeV) expert began a program at 8:40 AM that would fly a wire in the TeV every hour.

Operations put the Recycler into shot setup at 12:09 PM to stash antiprotons . (Any shot setup halts Pbar stacking and sending beam to Meson.)

Operations prepared Booster and the Preaccelerator for access at 12:57 PM.

Recycler stashing ended at 1:44 PM and the Booster access began.

The Booster work ended and everyone was out of the tunnel by 2:26 PM.

A Linac expert reported at 2:45 PM that he had finished tuning the I- Source .

At 3:01 PM, Operations reported that Booster was off due to a quadrupole power supply (B:QL2) problem.

The Booster power supply worked finished and Operations resumed sending beam to MiniBooNE at 4:28 PM. Pbar resumed stacking.

Operations established beam to MCenter at 4:37 PM.

A Booster RF station (BRF16) tripped off at 4:40 PM and then again at 6:07 PM. It reset okay both times.

Thursday April 14
The midnight shift began with Operations monitoring store 4097 and the Recycler stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 1:13 AM, BRF16 tripped off for the first of three times in three hours. It reset okay each time.

TeV experts began an end-of-store study at 7 AM.

Operations terminated store 4097 at 8:30 AM.

Operations put the TeV into shot setup at 8:47 AM.

At 10:44 AM, BRF16 tripped off. It reset okay.

Operations established combination store 4098 at 11:07 AM with an initial luminosity of 113.75E30. Pbar resumed stacking.

At 11:29 AM, a Pbar Debuncher power supply (D:H704) tripped off. Operators contacted an expert who replaced a SCR.

Pbar resumed stacking at 11:51 AM.

A Main Injector (MI) expert began a slip stacking study at 2:14 PM. She finished her study at 3:52 PM.

The Recycler began stashing antiprotons around 4 PM and ended at 5:06 PM. Pbar resumed stacking.

At 5:07 PM, MIRF2 tripped off three times. Three minutes later it tripped off repeatedly. An operator called an expert.

MCenter resumed taking beam 5:18 PM.

At 9:07 PM, MIRF14 tripped off for the first of four times in an hour. It reset okay each time.

At 9:16 PM, MI experts tried to conduct a study, but problems with a module made them cancel the study.

Friday April 15
The midnight shift began with Operations monitoring store 4098 and the stash, with Pbar stacking and conducting studies, and with MiniBooNE and MCenter taking beam.

At 1:12 AM, Operations reported a beam alarm at the AP1 transfer line. Operators investigated and called in an expert to help. Beam seemed to be mistuned.

At 2 AM, Operations reported that neither operators nor the expert were able to solve the AP1 line beam problem. Investigations continued.

At 4:22 AM, MI experts reported that a quadrupole in the P2 transfer line might be causing the problem. Another expert was called in. Operators continued to investigate.

At 6:02 AM, a MI expert reported that there was a focusing problem in the P1/P2 line.

The Plan for Friday and the Weekend The plan for today is to find the problem in the P1 or P2 transfer lines or to try and tune around the problem so stacking can resume. By 8 AM this morning, the investigation had centered on a possible bad quadrupole magnet. Investigations continue.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/13/2005   email Fermilab